Please help me. I have looked at this for an hour and I quit! Amy paid $76.49 for a pair of running shoes during a 15% off sale. What was the regular price? Thank you for any help give.
Answer by tutorcecilia(2144)

Original - (percent)(original = Discount [formula for price discount]
x-(.15x)=76.49 [plug-in the known values and use a variable (x) for the unknown]
x-.15x=76.49 [simplify; combine like-terms]
0.85x=76.49 [solve for the x-value]
x=76.59/0.85
x=89.99 [original price]
.
Check by plugging all of the values back into the original equation:
89.99-(.15)(89.99)=76.49
89.99-13.49=76.49
76.49-76.49 [checks out]